OpenStudy (anonymous):

Determine whether the series is convergent of divergent. sum of (4+3^n)/2^n from n=1 to infinite (I was told that this diverges because of the limit test but i dont understand why)

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ (4+3^n) }{ (2^n) }\]

OpenStudy (anonymous):

What is the value of the following limit: \[ \lim_{x \rightarrow \infty} \frac{4+3^x}{2^x}\]

OpenStudy (anonymous):

OH! when they said limit test i thought they meant the limit comparison test! So all i have to do is plug in infinite for n and it equals infinite

OpenStudy (anonymous):

eh... yeah. That's the general idea, though a mathematician would squirm if you said "plug in infinity and it equals infinity" :)

OpenStudy (anonymous):

haha i know :)